# Theory, and Design Guidelines

### v1

@@ Abstract
Optimization methods play a central role in signal processing, serving as the mathematical foundation for inference, estimation, and control. While classical iterative optimization algorithms provide interpretability and theoretical guarantees, they often rely on surrogate objectives, require careful hyperparameter tuning, and exhibit substantial computational latency. Conversely, machine learning (ML) offers powerful data-driven modeling capabilities but lacks the structure, transparency, and efficiency needed for optimization-driven inference. Deep unfolding has recently emerged as a compelling framework that bridges these two paradigms by systematically transforming iterative optimization algorithms into structured, trainable ML architectures. This article provides a tutorial-style overview of deep unfolding, presenting a unified perspective of methodologies for converting optimization solvers into ML models and highlighting their conceptual, theoretical, and practical implications. We review the foundations of optimization for inference and for learning, introduce four representative design paradigms for deep unfolding, and discuss the distinctive training schemes that arise from their iterative nature. Furthermore, we survey recent theoretical advances that establish convergence and generalization guarantees for unfolded optimizers, and provide comparative qualitative and empirical studies illustrating their relative trade-offs in complexity, interpretability, and robustness.
@@ I Introduction
In this article, we provide a tutorial-style overview of deep unfolding, combining a systematic taxonomy, a discussion of recent developments, and insights into the conceptual and theoretical advances that have shaped the field. Our goal is to provide an intuitive yet rigorous understanding of the capabilities and possibilities that arise when converting iterative optimization algorithms into ML architectures. We emphasize both the potential benefits of such hybrid designs, in terms of key measures such as interpretability, efficiency, and adaptability, as well as common challenges arising from their usage for inference in signal processing tasks. To that end, we begin by establishing the necessary preliminaries in optimization, focusing on the mathematical aspects that are most relevant for understanding the various forms of deep unfolding, and particularly on the formulations of optimization for inference and optimization for learning that serve as its conceptual foundation.
@@ II-A Optimization Problems
$\mathop{\rm minimize}\limits_{{\boldsymbol{s}}\in{\mathbb{R}}^{n}}\quad\mathcal{L}_{0}({\boldsymbol{s}}),\quad{\rm subject~to}\quad\mathcal{L}_{i}({\boldsymbol{s}})\leq 0,\quad i=1,\ldots,m,$
@@ II-B Iterative Optimization
For instance, consider an unconstrained optimization in (3), where $\mathcal{S}={\mathbb{R}}^{n}$, and the objective $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{s}};{\boldsymbol{x}})$ is convex and differentiable w.r.t. $\boldsymbol{s}$}. In this setting, one can iteratively update over ${\boldsymbol{s}}^{(k)}$, where $k=0,1,2,\ldots$ is the iteration index, and guarantee that under standard conditions the sequence $\{{\boldsymbol{s}}^{(k)}\}$ monotonically decreases the objective, i.e., $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{s}}^{(k+1)};{\boldsymbol{x}})<\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{s}}^{(k)};{\boldsymbol{x}})$, and converge to the global minimizer $\boldsymbol{s}$$^{*}$. The family of iterative optimizers that operate this way is referred to as descent methods. Descent methods that operate based on first-order derivatives of the objective, i.e., using gradients, are referred to as first-order methods. Arguably the most common first-order optimizer is gradient descent, which stems from the first-order multivariate Taylor series expansion of $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{s}};{\boldsymbol{x}})$ around the current iterate ${\boldsymbol{s}}^{(k)}$, to obtain an update equation
@@
In optimization problems where the objective function can be decomposed into the sum of two functions, $g(\cdot)+h(\cdot)$, with $g(\cdot)$ smooth and $h(\cdot)$ possibly nonsmooth, a widely-used descent method is based on proximal gradient descent. At each iteration $k$, this method updates ${\boldsymbol{s}}^{(k)}$ by locally approximating the smooth term $g(\cdot)$ via its second-order Taylor series expansion, where the Hessian matrix is approximated as $\frac{1}{\mu_{k}}{\boldsymbol{I}}$ for some step size $\mu_{k}\geq 0$. This representation leads to the iterative update rule
@@
Optimization-based inference rules, in which decisions are obtained as the solutions to optimization problems, exhibit several highly desirable properties that motivate their widespread use in signal processing and related fields:
@@
    Adaptability - A major advantage of iterative optimization-based inference is its inherent adaptability. Once an iterative solver is devised, it can typically be applied to different parameterizations of the objective (i.e., varying ${\boldsymbol{\theta}}^{\rm o}$) without redesigning the algorithmic structure. For example, the same ISTA solver can be used for sparse recovery problems under different measurement matrices $\boldsymbol{H}$}, where changing $\boldsymbol{H}$} merely modifies the instantiated objective $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{s}};{\boldsymbol{x}})$ while preserving the solver’s operational flow.
@@
    Scalability - Iterative optimizers are inherently scalable. Once an update rule is established, it can be applied seamlessly to problems of different dimensions, i.e., for varying signal sizes $n$ and measurement dimensions $m$. This dimension-agnostic nature facilitates the use of the same optimization-based inference mechanism across a broad range of problem instances without requiring any retraining or structural modification.
@@
    Surrogate Objectives - The objective guiding the inference rule is typically a surrogate of the true performance goal. For example, while the mathematical formulation of super-resolution as sparse recovery yields a tractable optimization problem, it may not fully capture the complex process of reconstructing high-resolution content from low-resolution, distorted measurements. Consequently, the optimal solution to the surrogate objective may not directly correspond to optimal task performance in practice.
@@ III Deep Unfolding
Figure 1: Illustration of considered methodologies for converting an iteration optimization-based inference mapping into an unfolded ML architecture. Trainable parameters are marked in red font.
@@ III-A Rationale
This rationale encompasses a remarkably broad family of learned optimizers that differ substantially in their specificity, degree of abstraction, and parameterization. This diversity arises from the dual parameterization inherent to optimization-based inference: the objective parameters ${\boldsymbol{\theta}}^{\rm o}$, which characterize the mathematical formulation of the task, and the hyperparameters ${\boldsymbol{\theta}}^{\rm h}$, which govern the behavior of the iterative solver. Depending on which of these elements are made learnable and how learning is integrated into the iterative process, different forms of deep unfolding emerge, each suited to distinct scenarios and offering unique benefits.
Specifically, one can identify four principal paradigms of deep unfolding:
@@
An illustration of the different architectural design approaches is depicted in Fig. 1. All methods result in the iterative optimizer with $K$ iterations being cast as an ML model. In the following subsections, we elaborate on each of these methodologies, providing their mathematical formulations, representative examples, and insights into their practical and theoretical gains, followed by a comparative discussion in Section V.
@@ III-B Learning Hyperparameters
This form of deep unfolding fully preserves the operation of the original iterative optimizer (thus satisfying P3), modifying only its hyperparameters, which are typically manually tuned, while preserving its adaptability to different objective parameters (P2). Its main motivation lies in enforcing an iterative optimizer to achieve its best possible performance under a fixed and predefined latency constraint (tackling C3), i.e., with a fixed number of iterations $K$. Unlike conventional methods for hyperparameter adaptation, such as backtracking or line search [29, Ch. 3], [30], learning the hyperparameters is carried out during an offline training phase, and thus, incurs no additional inference-time complexity.
@@ III-C Learning Objective Parameters
The second key advantage stems from the ability to cope with the inherent surrogate nature of optimization objectives (C1) and the frequent reliance on relaxed objectives for tractability (C2). For once, objective parameters, such as regularization coefficients in sparse recovery (e.g., the $\ell_{1}$ coefficient $\rho$ in ISTA), typically require manual tuning to achieve satisfactory results. Here, these parameters are instead learned from data, allowing the optimizer to automatically adjust its objective formulation. Furthermore, even when the original objective is an accurate representation of the task, the trainable nature of this methodology enables the design of alternative objectives that optimize for downstream performance measures. For example, it was shown in [31] that when optimization-based equalization is followed by symbol detection in quantized multiple-input multiple-output (MIMO) systems, learning the objective parameters allows the unfolded optimizer to adapt its criterion such that its outputs are most beneficial for the subsequent detection stage.
@@
While the gains discussed thus far relate primarily to conventional iterative optimization, learning objective parameters also brings forth notable advantages compared to black-box DNNs. Most importantly, this methodology retains full interpretability (P3): for any learned parameter set $\{{\boldsymbol{\theta}}^{\rm h}_{k},{\boldsymbol{\theta}}^{\rm o}_{k}\}$, one can explicitly identify the objective governing each iteration and the corresponding solver hyperparameters. This interpretability can be directly leveraged during training, as detailed in Subsection III-F. Moreover, the existence of an exact parameter configuration corresponding to the original non-learned optimizer provides a principled and effective initialization scheme. Initializing the unfolded model with the parameters of the original solver ensures that training begins from a well-understood, interpretable starting point and often leads to faster convergence and improved generalization.
@@ III-F Training Unfolded Architectures
The sequential structure of deep unfolded models provides an opportunity to incorporate losses associated with intermediate iterations. To formulate this mathematically, we focus on supervised learning as in (6) with sample-wise loss $l(\cdot)$ and without regularization (i.e., $\lambda=0$). By denoting the trainable parameters of the $k$th iteration as ${\boldsymbol{\theta}}_{k}$, i.e., ${\boldsymbol{\theta}}=\{{\boldsymbol{\theta}}_{k}\}$, and writing ${{\boldsymbol{s}}}^{(k)}({\boldsymbol{x}};{\boldsymbol{\theta}}_{1},\ldots,{\boldsymbol{\theta}}_{k})$ as the output of the $k$th iteration of an unfolded architecture with trainable parameters $\{{\boldsymbol{\theta}}_{k}\}$, we can identify three different training loss options:
@@
    End-to-end loss, which follows the common deep learning practice and evaluates the model based on its output, which here are $\mathcal{F}_{{\boldsymbol{\theta}}}({\boldsymbol{x}})={{\boldsymbol{s}}}^{(K)}({\boldsymbol{x}};{\boldsymbol{\theta}}_{1},\ldots,{\boldsymbol{\theta}}_{k})$, and thus
@@
    In (17b), $\{\alpha_{k}\}$ represent the weight assigned to each iteration, with a candidate setting proposed in [13] sets $\alpha_{k}=\log(1+k)$ to gradually assign more weight to latter iterations.

### v2

@@ Abstract
Optimization methods play a central role in signal processing, serving as the mathematical foundation for inference, estimation, and control. While classical iterative optimization algorithms provide interpretability and theoretical guarantees, they often rely on surrogate objectives, require careful hyperparameter tuning, and exhibit substantial computational latency. Conversely, machine learning (ML) offers powerful data-driven modeling capabilities but lacks the structure, transparency, and efficiency needed for optimization-driven inference. Deep unfolding has recently emerged as a compelling framework that bridges these two paradigms by systematically transforming iterative optimization algorithms into structured, trainable ML architectures. This article provides a tutorial-style overview of deep unfolding, presenting a unified perspective of methodologies for converting optimization solvers into ML models and highlighting their conceptual, theoretical, and practical implications. We review the foundations of optimization for inference and for learning, introduce four representative design paradigms for deep unfolding, and discuss the distinctive training schemes that arise from their iterative nature. To concretely illustrate the discussed principles, we use canonical examples of deep unfolding methodologies for representative problems in sparse recovery and communication optimization, while the presented methodology extends well beyond these specific settings. Furthermore, we survey recent theoretical advances that establish convergence and generalization guarantees for unfolded optimizers, and provide comparative qualitative and empirical studies illustrating their relative trade-offs in complexity, interpretability, and robustness.
@@ I Introduction
In this article, we provide a tutorial-style overview of deep unfolding, combining a systematic taxonomy, a discussion of recent developments, and insights into the conceptual and theoretical advances that have shaped the field. Our goal is to provide an intuitive yet rigorous understanding of the capabilities and possibilities that arise when converting iterative optimization algorithms into ML architectures. We emphasize both the potential benefits of such hybrid designs, in terms of key measures such as interpretability, efficiency, and adaptability, as well as common challenges arising from their usage for inference in signal processing tasks. We thus provide a tutorial-style treatment of deep unfolding that synthesizes recent methodological, theoretical, and practical developments into a unified and application-oriented framework, rather than a comprehensive survey of all related model-based learning paradigms. To that end, we begin by establishing the necessary preliminaries in optimization, focusing on the mathematical aspects that are most relevant for understanding the various forms of deep unfolding, and particularly on the formulations of optimization for inference and optimization for learning that serve as its conceptual foundation.
@@ II-A Optimization Problems
$\mathop{\rm minimize}\limits_{{\boldsymbol{s}}\in{\mathbb{R}}^{n}}\quad{\boldsymbol{L}}_{0}({\boldsymbol{s}}),\quad{\rm subject~to}\quad\mathcal{L}_{i}({\boldsymbol{s}})\leq 0,\quad i=1,\ldots,m,$
@@ II-B Iterative Optimization
For instance, consider an unconstrained optimization in (3), where $\mathcal{S}={\mathbb{R}}^{n}$, and the objective $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{s}};{\boldsymbol{x}})$ is convex and differentiable w.r.t. $\boldsymbol{s}$}. In this setting, one can iteratively update over ${\boldsymbol{s}}^{(k)}$, where $k=0,1,2,\ldots$ is the iteration index, and guarantee that under standard conditions the sequence $\{{\boldsymbol{s}}^{(k)}\}$ monotonically does not increase the objective, i.e., $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{s}}^{(k+1)};{\boldsymbol{x}})\leq\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{s}}^{(k)};{\boldsymbol{x}})$, and converges to a minimizer of the problem, or to the global minimizer $\boldsymbol{s}$$^{*}$ when the objective is convex. The family of iterative optimizers that operate this way is referred to as descent methods. Descent methods that operate based on first-order derivatives of the objective, i.e., using gradients, are referred to as first-order methods. Arguably the most common first-order optimizer is gradient descent, which stems from the first-order multivariate Taylor series expansion of $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{s}};{\boldsymbol{x}})$ around the current iterate ${\boldsymbol{s}}^{(k)}$, to obtain an update equation
@@
In optimization problems where the objective function can be decomposed into the sum of two functions, $g(\cdot)+h(\cdot)$, with $g(\cdot)$ smooth and $h(\cdot)$ possibly nonsmooth, a widely-used descent method is based on proximal gradient descent. At each iteration $k$, this method updates ${\boldsymbol{s}}^{(k)}$ by locally approximating the smooth term $g(\cdot)$ with a quadratic upper bound (which can be interpreted as a second-order Taylor approximation with a scaled identity Hessian surrogate), leading to the iterative update rule
@@
Optimization-based inference rules, in which decisions are obtained as the solutions to optimization problems, exhibit several desirable properties that motivate their widespread use in signal processing and related fields:
@@
    Adaptability - A major advantage of iterative optimization-based inference is its inherent adaptability. Once an iterative solver is devised, it can typically be applied to different parameterizations of the objective (i.e., varying ${\boldsymbol{\theta}}^{\rm o}$) without redesigning the algorithmic structure. For example, the same ISTA solver can be used for sparse recovery under different measurement matrices $\boldsymbol{H}$}, where changing $\boldsymbol{H}$} merely modifies the instantiated objective $\mathcal{L}_{{\boldsymbol{\theta}}^{\rm o}}({\boldsymbol{s}};{\boldsymbol{x}})$ while preserving the solver’s operational flow.
@@
    Scalability - Iterative optimizers are inherently scalable. Once an update rule is established, it can be applied seamlessly to problems of different dimensions, i.e., for varying signal sizes $n$ and measurement dimensions $m$. This dimension-agnostic nature facilitates the use of the same optimization-based inference mechanism across a broad range of problem instances without requiring any retraining or structural modification. While scalability is not unique to optimization-based inference (e.g., convolutional or graph neural networks can also operate across varying spatial or graph dimensions), optimization-derived update rules are often inherently dimension-agnostic, allowing the same solver structure to be directly applied to problem instances of different sizes without architectural modification.
@@
    Surrogate Objectives - The objective guiding the inference rule is typically a surrogate of the true performance goal. For example, while the mathematical formulation of super-resolution as sparse recovery yields a tractable optimization problem, it may not fully capture the complex process of reconstructing high-resolution content from low-resolution, distorted measurements. Consequently, the solution to the surrogate objective may not directly correspond to optimal task performance in practice.
@@ III Deep Unfolding
Figure 1: Illustration of considered methodologies for converting an iterative optimization-based inference mapping into an unfolded ML architecture. Trainable parameters are marked in red font.
@@ III-A Rationale
Deep unfolding is part of a broader family of approaches often referred to as learn-to-optimize [29, 30], which combine ML with iterative optimization algorithms. Beyond finite-depth unrolling, this family includes implicit-layer approaches such as deep equilibrium models [31], as well as plug-and-play methods [32] and related algorithm-inspired architectures [10] that embed learned modules within iterative optimization frameworks. The distinguishing feature of deep unfolding is the explicit truncation of an iterative solver into a fixed-depth architecture trained end-to-end, whereas implicit or plug-and-play paradigms typically preserve the convergence-oriented viewpoint and differentiate through fixed-point conditions or outer optimization loops. In this article, we focus specifically on the deep unfolding paradigm and its unique properties.
The guiding rationale of deep unfolding encompasses a remarkably broad family of learned optimizers that differ substantially in their specificity, degree of abstraction, and parameterization. This diversity arises from the dual parameterization inherent to optimization-based inference: the objective parameters ${\boldsymbol{\theta}}^{\rm o}$, which characterize the mathematical formulation of the task, and the hyperparameters ${\boldsymbol{\theta}}^{\rm h}$, which govern the behavior of the iterative solver. Depending on which of these elements are made learnable and how learning is integrated into the iterative process, different forms of deep unfolding emerge, each suited to distinct scenarios and offering unique benefits. The selection of the underlying iterative solver is a design choice on its own. A given optimization problem can often be addressed by multiple algorithms, each inducing a distinct computational structure and parameterization. When unfolded, these structural differences directly affect the resulting architecture, including the number of learnable hyperparameters and the degree of abstractness. Accordingly, the choice of solver is not merely an optimization detail, but a fundamental architectural decision in deep unfolding design.
For a given iterative optimizer, one can identify four principal paradigms of deep unfolding:
@@
An illustration of the different architectural design approaches is depicted in Fig. 1. All methods result in the iterative optimizer with $K$ iterations being cast as an ML model. In the following subsections, we elaborate on each of these methodologies, providing their mathematical formulations, representative examples, and insights into their practical and theoretical gains, followed by a comparative discussion in Section V. While, for clarity of exposition, the illustrative examples in the sequel focus on optimization problems written in regularized (unconstrained) form, the deep unfolding methodologies described in this section apply equally to minimax [15] or to explicitly constrained formulations (as in (1)) [33, 34], by unfolding the corresponding iterative solvers.
@@ III-B Learning Hyperparameters
This form of deep unfolding fully preserves the operation of the original iterative optimizer (thus satisfying P3), modifying only its hyperparameters, which are typically manually tuned, while preserving its adaptability to different objective parameters (P2). Its main motivation lies in enforcing an iterative optimizer to achieve its best possible performance under a fixed and predefined latency constraint (tackling C3), i.e., with a fixed number of iterations $K$. Unlike conventional methods for hyperparameter adaptation, such as backtracking or line search [35, Ch. 3], [36], learning the hyperparameters is carried out during an offline training phase, and thus, incurs no additional inference-time complexity.
@@ III-C Learning Objective Parameters
The second key advantage stems from the flexibility to adapt surrogate (C1) or relaxed objectives (C2) to the data. While learning objective parameters does not eliminate the surrogate modeling assumption, it enables iteration-dependent adaptation of the objective formulation. For example, objective parameters, such as regularization coefficients in sparse recovery (e.g., the $\ell_{1}$ coefficient $\rho$ in ISTA), typically require manual tuning to achieve satisfactory results. In LISTA-type architectures, different layers effectively employ different regularization strengths, allowing the unfolded optimizer to adjust the surrogate trade-off across iterations. This adaptive mechanism can empirically mitigate mismatches between the surrogate objective and the underlying data distribution, and can further tailor the optimization criterion to downstream performance measures. Furthermore, even when the original objective is an accurate representation of the task, the trainable nature of this methodology enables the design of alternative objectives that optimize for downstream performance measures. For example, it was shown in [37] that when optimization-based equalization is followed by symbol detection in quantized multiple-input multiple-output (MIMO) systems, learning the objective parameters allows the unfolded optimizer to adapt its criterion such that its outputs are most beneficial for the subsequent detection stage.
@@
While the gains discussed thus far relate primarily to conventional iterative optimization, learning objective parameters also brings forth notable advantages compared to black-box DNNs. Learning objective parameters preserves the algorithmic structure of the unfolded architecture, allowing one to explicitly identify the objective and solver hyperparameters associated with each iteration. This brings forth some level of interpretability (P3) which is structural here rather than global. In particular, when objectives vary across layers without explicit constraints tying them to a single underlying problem, the resulting network executes a sequence of optimization-inspired transformations. Moreover, the existence of an exact parameter configuration corresponding to the original non-learned optimizer provides a principled and effective initialization scheme. Initializing the unfolded model with the parameters of the original solver ensures that training begins from a well-understood, interpretable starting point and often leads to faster convergence and improved generalization.
Finally, recent works have shown that in some cases, the learning objective parameters can give rise to architectures resembling modern transformer networks. For instance, [38] derives self-attention-like operations by unrolling an optimization problem associated with sparse rate reduction, where the parameters of the underlying representation model are learned. Similarly, [39] obtains lightweight transformer architectures by unrolling a graph-based signal interpolation problem, in which the graph structure (i.e., edge weights) is learned from data. These examples highlight how learning the parameters of the optimization objective can induce rich, interpretable neural modules.
@@ III-F Training Unfolded Architectures
Parameter Tying
The different formulations presented in the previous subsections specify how one can convert an iteration of an iterative optimizer into a trainable parameterized ML model. In learning these parameters $\boldsymbol{\theta}$}, one can choose whether to tie parameters across iterations or allow iteration-dependent parameterizations. Tying weights preserves a closer correspondence with the original iterative algorithm, reduces the number of trainable parameters, and often enhances stability and generalization. Alternatively, allowing iteration-specific parameters increases expressivity at the cost of a larger model. Both regimes are fully compatible with the deep unfolding paradigm, and the choice between them reflects a trade-off between model capacity and structural regularization.
@@
Clearly, unfolded networks can also be trained using alternative ML self-supervised criteria, such as data-consistency-based losses or Noise2Noise-type formulations [48]. The key advantage of deep unfolding is that, in addition to these generic learning strategies, it naturally provides a principled objective function that can serve as a sample-wise training loss.
@@
The sequential structure of deep unfolded models provides an opportunity to incorporate losses associated with intermediate iterations. To formulate this mathematically, we focus on supervised learning as in (6) with sample-wise loss $l(\cdot)$ and without regularization (i.e., $\lambda=0$). By denoting the trainable parameters of the $k$th iteration as ${\boldsymbol{\theta}}_{k}$ (focusing on the generic case where the parameters can change between iteration), i.e., ${\boldsymbol{\theta}}=\{{\boldsymbol{\theta}}_{k}\}$, and writing ${{\boldsymbol{s}}}^{(k)}({\boldsymbol{x}};{\boldsymbol{\theta}}_{1},\ldots,{\boldsymbol{\theta}}_{k})$ as the output of the $k$th iteration of an unfolded architecture with trainable parameters $\{{\boldsymbol{\theta}}_{k}\}$, we can identify three different training loss options:
@@
    End-to-end loss, which follows the common deep learning practice and evaluates the model based on its output, which here is $\mathcal{F}_{{\boldsymbol{\theta}}}({\boldsymbol{x}})={{\boldsymbol{s}}}^{(K)}({\boldsymbol{x}};{\boldsymbol{\theta}}_{1},\ldots,{\boldsymbol{\theta}}_{k})$, and thus
@@
    In (17b), $\{\alpha_{k}\}$ represent the weight assigned to each iteration, with a candidate setting proposed in [13] sets $\alpha_{k}=\log(1+k)$ to gradually assign more weight to later iterations.